# On the nullspace of split graphs

### v2

@@ Abstract
We study the nullspace of the adjacency matrix of split graphs, whose vertex set can be partitioned into a clique and an independent set. We introduce the clique-kernel, a subspace that decides whether clique vertices lie in the support of a kernel eigenvector, and we prove that its dimension is at most one. This yields the the formula
@@
which fully describes the nullity of a split graph in terms of the biadjacency submatrix $R$. We also analyze unbalanced split graphs through the concept of swing vertices and characterize the structure of their kernel supports. Furthermore, we study the behavior of the nullspace under Tyshkevich composition and derive a closed formula for the determinant. These results provide a unified algebraic framework for understanding when a split graph is singular and how its combinatorial structure determines its nullspace.
@@
Split graph, Nullspace, Graph Kernel, Adjacency Matrix, Tyshkevich Composition, Determinant
MSC:
@@ 1 Introduction
    An explicit determinant formula for split graphs (Section 8) expressed in terms of the matrix $R$ using the Schur complement and the matrix determinant lemma, which facilitates computational verification of singularity.
@@
Throughout this work, we demonstrate that the nullspace of split graphs exhibits rich structural properties that can be understood through the interplay between the clique-independent set s-partition and the rank properties of the biadjacency submatrix $R$. Our results provide both theoretical insights into the spectral theory of structured graph classes and computational tools for determining singularity and nullity of split graphs.
@@ 2 Notation
If $G$ is a graph, then $N_{G}(v)=\{u\in V(G):uv\in E(G)\}$ is the neighborhood of the vertex $v$ in $G$ (we can write $N(v)$ if $G$ is clear from the context). The symbols $\omega(G)$ and $\alpha(G)$ denote, respectively, the clique number and the independence number of $G$.
A graph $G\neq(\varnothing,\varnothing)$ is said to be split if there exists a clique $K$ in $G$ and an independent set $S$ in $G$ such that $V(G)=K\dot{\cup}S$. In such a case, we use the notation $Sp$ instead of $G$ and we say that the pair $(K,S)$ is an s-partition for $Sp$. Directly from the definition, it follows that every induced subgraph of a split graph is also split. We call clique vertices the elements of $K$, and independent vertices the elements of $S$. Two s-partitions $(K,S)$ and $(K^{\prime},S^{\prime})$ for $Sp$ are equal if $K=K^{\prime}$ and $S=S^{\prime}$.
Figure 1: Split graph SpSp (clique vertices in black, independent in white).
@@ 3 Nullspace of split graphs
When studying split graphs one is naturally led to investigate the column space of the biadjacency matrix $R$ and its interplay with the all-ones vector $\operatorname{1\!\!1}$. The next lemma establishes a general matrix-theoretic criterion that is pivotal for such analysis. It characterizes the membership of a vector in the image of a matrix via the containment of nullspaces, a formulation that proves especially useful when applied to rank-one matrices like $J=\operatorname{1\!\!1}\operatorname{1\!\!1}^{t}$.
@@
    By Lemma 3.1, we have $\operatorname{\mathcal{N}}(R^{t})\subset\operatorname{\mathcal{N}}(J)$. Now, for any $(x_{K},x_{S})\in\operatorname{\mathcal{N}}(Sp)$, by (1) we have $R^{t}x_{K}=0$, so $x_{K}\in\operatorname{\mathcal{N}}(R^{t})\subset\operatorname{\mathcal{N}}(J)$, meaning $\operatorname{1\!\!1}^{t}x_{K}=0$. Then by part (4), $x_{K}=0$. Therefore, $\operatorname{Supp}(Sp)\subset S$.∎
This theorem provides fundamental tools for understanding the nullspace structure of split graphs. In particular, item (5) gives a sufficient condition for when all support vertices lie in the independent set $S$, which has important combinatorial implications.
It is important to note that the converse of item (5) of Theorem 3.2 is not true. In other words, there exist split graphs with $\operatorname{Supp}(S)\subset S$ and $\operatorname{1\!\!1}\notin\operatorname{\mathcal{R}}(R)$. To see that, consider the split graph $Sp(K,S)$ in Figure 2.
@@
Direct computations shows that $\operatorname{Supp}(Sp)=\{5,\ldots,9\}=S$. However, the equation $Rx=\operatorname{1\!\!1}$ has no solution because
@@
$\operatorname{Cker}{\left(Sp\right)}=\operatorname{\mathcal{N}}(R^{t})\cap\operatorname{\mathcal{R}}((I-J)^{-1}R)\subset\mathbb{R}^{|K|}.$
@@
The relationship between the sizes of the clique and independent set in a split graph has implications for its singularity. The following result establishes simple combinatorial criteria that guarantee when a split graph must be singular based solely on the cardinalities of its s-partition. In particular, we show that an imbalance in favor of the independent set always forces singularity, while in the balanced case, the singularity of the entire graph reduces to the singularity of the biadjacency matrix. This provides a bridge between the graph’s global spectral properties and the local structure encoded in the matrix $R$.
@@
Regularity conditions within the clique or independent set impose strong constraints on the nullspace structure of a split graph. This theorem reveals how degree homogeneity among clique vertices confines the support entirely to the independent set, while degree homogeneity among independent vertices forces a specific linear dependence in the kernel of $R$. These results demonstrate how local symmetry conditions—all vertices in one part having identical degrees—propagate to global spectral properties, offering combinatorial tools to control the location of support vertices and the structure of the nullspace.
@@
    which shows that $\operatorname{1\!\!1}\in\operatorname{\mathcal{R}}(R^{t})$. By Lemma 3.1, this implies $\operatorname{\mathcal{N}}(R)\subset\operatorname{\mathcal{N}}(J)$. ∎
@@ 4 Unbalanced split graphs
We say that a vertex $w$ of a split graph $Sp$ is swing if there exist s-partitions $(K,S),(K^{\prime},S^{\prime})$ for $Sp$ such that $w\in K$ and $w\in S^{\prime}$. Let $W(Sp)$ be the set of all swing vertices of $Sp$. Then, it follows that $Sp$ is balanced if and only if $W(Sp)=\varnothing$. However, $Sp\setminus W(Sp)$ may not be balanced in general. It is easy to see that $w\in W(Sp)$ if and only if there are s-partitions $(K,S),(K^{\prime},S^{\prime})$ for $Sp$ such that $w\in K$, $w\in S^{\prime}$ and
$K\setminus\{w\}=N_{Sp}(w)=K^{\prime},$ (1)
$(|K|,|S|)=(\omega(Sp),\alpha(Sp)-1),\ (|K^{\prime}|,|S^{\prime}|)=(\omega(Sp)-1,\alpha(Sp)).$
If $Sp$ is a split graph, let:
@@
Then, by the definition of swing vertex, it is clear that
$V(Sp)=K^{*}(Sp)\ \dot{\cup}\ W(Sp)\ \dot{\cup}\ S^{*}(Sp).$ (2)
@@
Although (3) agrees with (1), we think there is an issue with (3): it depends on the s-partition. To see that, consider the split graph $Sp=abc\approx P_{3}$ with the s-partitions $(K,S)=(\{a,b\},c)$ and $(K^{\prime},S^{\prime})=(a,\{b,c\})$. By our definition, $W(Sp)=\{a,c\}$, which is an invariant of $Sp$. However, using (3), we observe that: 1) in $Sp(K,S)$, $a$ is swing but $c$ is not; 2) in $Sp(K^{\prime},S^{\prime})$, $c$ is swing but $a$ is not. For this reason, we decided to modify (3). As we have seen above, this choice allows us to immediately obtain the useful vertex-decomposition (2).
Next theorem is a translation to our setting of a known result that characterizes the s-partitions of an unbalanced split graph.
Theorem 4.1 (Cheng, Collins, Trenk, [3], page 2348).
@@
    If $W=\{w\}$, then
    $\mathcal{KS}=\{(K^{*}\cup w,S^{*}),(K^{*},S^{*}\cup w)\};$
    moreover, $K^{*}\cup w$ is the unique maximum clique of $\mathcal{K}(Sp)$ and $S^{*}\cup w$ is the unique maximum independent set of $\mathcal{S}(Sp)$.
@@
    If $W$ is a clique of size $\geq 2$, then
    $\mathcal{KS}=\{(K^{*}\cup W,S^{*})\}\cup\{(K^{*}\cup(W\setminus\{w\}),S^{*}\cup w):w\in W\};$
    moreover, $K^{*}\cup W$ is the unique maximum clique of $\mathcal{K}(Sp)$ and $|S^{*}\cup w|=\alpha(Sp)$.
@@
    If $W$ is an independent set of size $\geq 2$, then
    $\mathcal{KS}=\{(K^{*},S^{*}\cup W)\}\cup\{(K^{*}\cup w,S^{*}\cup(W\setminus\{w\})):w\in W\};$
    moreover, $S^{*}\cup W$ is the unique maximum independent set of $\mathcal{S}(Sp)$ and $|K^{*}\cup w|=\omega(Sp)$.
The following proposition is a refinement of Theorem 4.1, because, essentially, tells us that $\mathcal{K}(Sp)$” and $\mathcal{S}(Sp)$” can be replaced with “SpSp” in Theorem 4.1.
Theorem 4.2.
@@
    Let $K$ be a maximum clique of $Sp$. Since $Sp$ is unbalanced, we know that $\omega(Sp)+\alpha(Sp)=|Sp|+1$, and so $\alpha-1=|Sp|-\omega=|V(Sp)\setminus K|$. Let $S$ be a maximum independent set of $Sp$. Since $|K\cap S|\leq 1,|S|=\alpha$ and $|V(Sp)\setminus K|=\alpha-1$, it must be $S=(V(Sp)\setminus K)\dot{\cup}w$, for some $w\in K$. Consequently, $V(Sp)\setminus K$ is an independent set of $Sp$. Thus, $(K,V(Sp)\setminus K)$ is an s-partition for $Sp$, which means that $K\in\mathcal{K}(Sp)$.
@@
    Let $S$ be a maximum independent set of $Sp$. Then, $S$ is a maximum clique of the graph $\overline{Sp}$ (the complement of $Sp$), which is clearly split and unbalanced. From (1), we know that $S\in\mathcal{K}(\overline{Sp})$, which means that $V(\overline{Sp})\setminus S$ is an independent set of $\overline{Sp}$. Hence, $V(Sp)\setminus S$ is a clique of $Sp$, which shows that $S\in\mathcal{S}(Sp)$.
Surprisingly, Theorem 4.2 does not hold in general for balanced split graph. In fact, every balanced $Sp(K,S)$ with an independent vertex of degree |K|1|K|-1 is a counterexample ($P_{4}$ is the “simplest” one). In other words, a balanced split graph $Sp(K,S)$ has always a unique s-partition, but $K$ might not be the only maximum clique of $Sp$. An analogous phenomenon can be observed for maximum independent sets. The following result clarifies this behavior of the balanced split graphs.
@@
    If $x$ is a vertex of $Sp$ of degree |K|1|K|-1, then $x$ must be independent (otherwise, $x$ would be swing). Now, observe that $N(x)\cup x$ is a maximum clique of $Sp$, different from $K$.
    For the converse, suppose that $K^{\prime}$ is a maximum clique of $Sp$, different from $K$. Since $|K^{\prime}\cap S|\leq 1$, it must be $K^{\prime}=(K\setminus\{w\})\dot{\cup}\{x\}$, for some $w\in K$ and $x\in S$. Since $x\in K^{\prime}$, we have $K\setminus\{w\}\subset N(x)$. Since $\in S, xx$ is not adjacent to any vertex of $S$, which means that $N(x)\subset K$. If $w\in N(x)$, then $N(x)=K$ and hence $x\in W(Sp)=\varnothing$. This absurdity implies $N(x)=K\setminus\{w\}$, i.e., $\deg(x)=|K|-1$.
@@
    If $x$ is a vertex of $Sp$ such that $N(x)\cap S=\{w\}$ for some $w$, it is clear that $x\in K$ and $(I\setminus\{w\})\dot{\cup}\{x\}$ is a maximum independent set different from $S$.
    For the converse, suppose that $S^{\prime}$ is a maximum independent set of $Sp$, different from $S$. Since $|K\cap S^{\prime}|\leq 1$, it must be $S^{\prime}=(S\setminus\{w\})\dot{\cup}\{x\}$, for some $w\in S$ and $x\in K$. Since $x\in K$, $K\setminus x\subset N(x)$. Since $\in S^{\prime}, xx$ is not adjacent to any vertex of $S\setminus\{w\}$. Notice that $w\in N(x)$ (otherwise, $x$ would be swing). Therefore, $N(x)=(K\setminus x)\dot{\cup}\{w\}$, which means that $|N(x)\cap S|=1$.
The following theorem states that in unbalanced split graphs, supported vertices must lie in the intersection of all the maximum independent sets. This has significant implications for understanding the nullspace structure of unbalanced split graphs.
@@
Consider an arbitrary s-partition $(K,S)$ for $Sp$ such that $|S|=\alpha(Sp)$. The existence of such is guaranteed by Theorem 4.1. Then, by Theorem 4.1, there exists a vertex $w\in S$ such that $N(w)=K$, which means that $R_{*w}=\operatorname{1\!\!1}$. Since $\operatorname{1\!\!1}\in\operatorname{\mathcal{R}}(R)$, we conclude by Theorem 3.2(5) that $\operatorname{Supp}(Sp)\subset S$. Therefore, $\operatorname{Supp}(Sp)$ is contained in the intersection of all the maximum independent sets in $\mathcal{S}(Sp)$. Now, use Theorem 4.2(2). ∎
@@
If $Sp$ is threshold, then $\operatorname{Supp}(Sp)$ is contained in the intersection of all the maximum independent sets of $Sp$.
@@
Follows from the previous discussion and from Theorem 4.4. ∎
Two vertices $u,v$ in a graph $G$ are said to be twins in $G$ if
$N_{G}(u)\setminus v=N_{G}(v)\setminus u.$
It is easy to check that “to be twins” is an equivalence relation over $V(G)$, and the equivalence classes $[v]$ are cliques or independent sets of $G$.
Lemma 4.6.
Let $Sp$ be a split graph. If $w\in W(Sp)$, then $[w]=W(Sp)$.
Proof.
If $W(Sp)=\{w\}$, the inclusion $W(Sp)\subset[w]$ is obvious. If $|W(Sp)|\geq 2$, let $w^{\prime}\in W(Sp)\setminus\{w\}$. Since $w$ and $w^{\prime}$ are swing, there are s-partitions $(K,S),(K^{\prime},S^{\prime})$ for $Sp$ such that
$w\in S,\ w^{\prime}\in S^{\prime},\ |S|=\alpha(Sp)=|S^{\prime}|,\ N(w)=K,\text{ and }\ N(w^{\prime})=K^{\prime}.$
Suppose that $K\neq K^{\prime}$. If $x\in K\setminus K^{\prime}$, then $x\in S^{\prime}$. Hence, $x\in W(Sp)$, and so $K^{\prime}=N(x)=K\setminus\{x\}$. Now, observe that $S^{\prime}=S\dot{\cup}\{x\}$: this is impossible because $S$ and $S^{\prime}$ are maximum independent sets. Therefore, $K=K^{\prime}$, which means that $w^{\prime}$ and $w$ are twins. We have just proven the inclusion $W(Sp)\subset[w]$.
To prove the reverse inclusion, consider a vertex $w^{\prime}$ with $N(w^{\prime})\setminus\{w\}=N(w)\setminus\{w^{\prime}\}$. Since $w$ is swing, there is an s-partition $(K,S)$ for $Sp$ such that $w\in S$ and $N(w)=K$. Then, $N(w^{\prime})=K\setminus\{w^{\prime}\}$, which means that $w^{\prime}\in W(Sp)$. Hence, $w]\subset W(Sp)$. ∎

### v3

@@ Abstract
We study the nullspace of the adjacency matrix of split graphs, whose vertex set can be partitioned into a clique and an independent set. We introduce the clique-kernel, a subspace that decides whether clique vertices lie in the support of a kernel eigenvector, and we prove that its dimension is at most one. This yields the formula
@@
which fully describes the nullity of a split graph in terms of its biadjacency submatrix $R$. We also analyze unbalanced split graphs through the concept of swing vertices and characterize the structure of their kernel supports. Furthermore, we study the behavior of the nullspace under Tyshkevich composition and derive a closed formula for the determinant.
@@
Split graph, nullspace, adjacency matrix, Tyshkevich composition, determinant
2020 MSC:
@@ 1 Introduction
    An explicit determinant formula for split graphs (Section 8) expressed in terms of the matrix $R$ using the Schur complement and Cauchy’s formula, which facilitates computational verification of singularity.
@@ 2 Notation
If $G$ is a graph, then $N_{G}(v)=\{u\in V(G):uv\in E(G)\}$ is the neighborhood of the vertex $v$ in $G$. The degree in $G$ of a vertex $v$ (i.e., $|N_{G}(v)|$) is denoted by $\deg_{G}(v)$. The symbols $\omega(G)$ and $\alpha(G)$ denote, respectively, the clique number and the independence number of $G$. When $G$ is clear from the context, we can write $N(v),\deg(v),\omega$ and $\alpha$ instead of $N_{G}(v),\deg_{G}(v),\omega(G)$ and $\alpha(G)$, respectively.
[…]
Figure 1: Split graph (clique vertices in black, independent in white).
@@ 3 Nullspace of split graphs
When studying split graphs one is naturally led to investigate the column space of the biadjacency matrix $R$ and its interplay with the all-ones vector $\operatorname{1\!\!1}$. The next lemma establishes a general matrix-theoretic criterion that is pivotal for such analysis.
@@
    By Section 3, we have $\operatorname{\mathcal{N}}(R^{t})\subset\operatorname{\mathcal{N}}(J)$. Now, for any $(x_{K},x_{S})\in\operatorname{\mathcal{N}}(Sp)$, by (1) we have $R^{t}x_{K}=0$, so $x_{K}\in\operatorname{\mathcal{N}}(R^{t})\subset\operatorname{\mathcal{N}}(J)$, meaning $\operatorname{1\!\!1}^{t}x_{K}=0$. Then by part (4), $x_{K}=0$. Therefore, $\operatorname{Supp}(Sp)\subset S$.∎
This theorem provides tools for understanding the nullspace structure of split graphs. In particular, item (5) gives a sufficient condition for when all support vertices lie in the independent set $S$.
It is important to note that the converse of item (5) of Theorem 3.2 is not true. In other words, there exist split graphs with $\operatorname{Supp}(Sp(K,S))\subset S$ and $\operatorname{1\!\!1}\notin\operatorname{\mathcal{R}}(R)$. To see that, consider the split graph $Sp(K,S)$ in Figure 2.
@@
Direct computations show that $\operatorname{Supp}(Sp)=\{5,\ldots,9\}=S$. However, the equation $Rx=\operatorname{1\!\!1}$ has no solution, otherwise we end up with the absurdity
@@
$\operatorname{Cker}{\left(Sp\right)}:=\operatorname{\mathcal{N}}(R^{t})\cap\operatorname{\mathcal{R}}((I-J)^{-1}R)\subset\mathbb{R}^{|K|}.$
@@
The relationship between the sizes of the clique and independent set in a split graph has implications for its singularity. The following result establishes simple combinatorial criteria that guarantee when a split graph must be singular based solely on the cardinalities of its s-partition. In particular, we show that an imbalance in favor of the independent set always forces singularity, while in the balanced case, the singularity of the entire graph reduces to the singularity of the biadjacency matrix.
@@
Regularity conditions within the clique or independent set impose strong constraints on the nullspace structure of a split graph. This theorem reveals how degree homogeneity among clique vertices confines the support entirely to the independent set, while degree homogeneity among independent vertices forces a specific linear dependence in the kernel of $R$.
@@
    which shows that $\operatorname{1\!\!1}\in\operatorname{\mathcal{R}}(R^{t})$. By Section 3, this implies $\operatorname{\mathcal{N}}(R)\subset\operatorname{\mathcal{N}}(J)$. ∎
@@ 4 Unbalanced split graphs
We say that a vertex $w$ of a split graph $Sp$ is swing in $Sp$ if there exist s-partitions $(K,S),(K^{\prime},S^{\prime})$ for $Sp$ such that $w\in K$ and $w\in S^{\prime}$. Let $W(Sp)$ be the set of all swing vertices of $Sp$. Clearly, $W(Sp)=W(\overline{Sp})$, and $Sp$ is balanced if and only if $W(Sp)=\varnothing$. However, $Sp\setminus W(Sp)$ may not be balanced in general (take $Sp\approx P_{3}$). If $Sp$ is a split graph, let:
@@
Then, by the above definitions, it is clear that
$V(Sp)=W(Sp)\ \dot{\cup}\ K^{*}(Sp)\ \dot{\cup}\ S^{*}(Sp).$ (1)
@@
There is an issue with (2): it depends on the s-partition. To see that, consider the split graph $Sp=abc\approx P_{3}$ with the s-partitions $(K,S)=(\{a,b\},c)$ and $(K^{\prime},S^{\prime})=(\{b,c\},a)$. By our definition, $W(Sp)=\{a,c\}$, which is an invariant of $Sp$. However, using (2), we observe that: 1) in $Sp(K,S)$, $a$ is swing but $c$ is not; 2) in $Sp(K^{\prime},S^{\prime})$, $c$ is swing but $a$ is not. For this reason, we decided to modify (2). As we have seen above, this choice allows us to immediately obtain the useful vertex-decomposition (1).
A graph $G$ is called an NG-graph if $\chi(G)+\chi(\overline{G})=|G|+1$, where $\chi(G)$ is the chromatic number of $G$. In particular, a split graph is NG if and only if it is unbalanced, because $\chi(G)=\omega(G)$ when $G$ is split, and $\alpha(Sp)+\omega(Sp)=|Sp|+1$ when $Sp$ is unbalanced. In [3] the authors study some structural and combinatorial properties of unbalanced split graphs in terms of what they call the $ABC$-partition of $G$, where:
$A_{G}=\{v\in V(G):\deg_{G}(v)=\chi(G)-1\},$
$B_{G}=\{v\in V(G):\deg_{G}(v)>\chi(G)-1\},$
$C_{G}=\{v\in V(G):\deg_{G}(v)<\chi(G)-1\}.$
The “GG” sub-index will be omitted if $G$ is clear from the context.
As we will see soon in Theorem 4.4, $ABC$-partition and $WK^{*}S^{*}$-partition (1) are essentially the same vertex decomposition. This will take several steps: Sections 4, 4 and 4. As a consequence of Theorem 4.4, we obtain a translation (Section 4), in terms of $WK^{*}S^{*}$-partition, of Theorem 10 in [3], which characterizes the s-partitions of an unbalanced split graph in terms of its $ABC$-partition. Moreover, thanks to Section 4, we obtain a slightly stronger version of the cited result. Section 4 is fundamental to subsequently understand the support of unbalanced split graphs.
Lemma 4.1.
If $Sp(K,S)$ is unbalanced, then $A_{Sp}\subset W(Sp)$.
Proof.
First of all, recall that $(|K|,|S|)\in\{(\omega,\alpha-1),(\omega-1,\alpha)\}$, since $Sp$ is unbalanced. Let $x\in A$. We have two cases: 1) $x\in K$; 2) $x\in S$.
  1. (1).
    If $x\in K$, then $K\setminus x\subset N(x)$.
    If $|K|=\omega-1$, then $N(x)=(K\setminus x)\dot{\cup}y$, for some $y\in S$. Since $N(y)\subset K$, we have $\deg(y)\leq\omega-1$. If $\deg(y)<\omega-1$, then $y\in C$, contradicting that there are no edges between $A$ and $C$ (see item 5 of Theorem 2 in [3]. So, $\deg(y)=\omega-1$, and hence $N(y)=K$. Therefore, $x\in W$ because $(N(x),(S\setminus y)\dot{\cup}x)$ is an s-partition for $Sp$ where $x$ is independent.
    If $|K|=\omega$, then $N(x)=K\setminus x$, which means that $x$ has no neighbors in $S$. Now, it is easy to check that $(N(x),S\dot{\cup}x)$ is an s-partition for $Sp$ where $x$ is independent.
  2. (2).
    If $x\in S$, then $x$ is a clique vertex in $\overline{Sp}(S,K)$ with
    $\deg_{\overline{Sp}}(x)=|Sp|-1-\deg(x)=|Sp|-\omega=\alpha-1=\omega(\overline{Sp})-1$
    (recall that $\alpha+\omega=|Sp|+1$, because $Sp$ is unbalanced), which means that $x\in A_{\overline{Sp}}$. Then, we can apply (1) to infer that $x\in W(\overline{Sp})=W$. ∎
Lemma 4.2.
If $Sp$ is unbalanced, then $B_{Sp}\subset K^{*}(Sp)$.
Proof.
Assume that $x\in B$ but there exists an s-partition $(K,S)$ for $Sp$ such that $x\in S$. Since $N(x)\subset K$ and $\deg(x)>\omega-1$, we have $\deg(x)=\omega$. Thus, $N(x)=K$ and $|K|=\omega$. Now, observe that $K\dot{\cup}x$ is a clique of size $\omega+1$ in $Sp$, which is absurd. ∎
Lemma 4.3.
If $Sp$ is unbalanced, then $C_{Sp}\subset S^{*}(Sp)$.
Proof.
Assume that $x\in C$ but there exists an s-partition $(K,S)$ for $Sp$ such that $x\in K$. Since $K-x\subset N(x)$ and $\deg(x)<\omega-1$, we have $\deg(x)=\omega-2$. Thus, $N(x)=K\setminus x$ and $|K|=\omega-1$. Now, observe that $S\dot{\cup}x$ is an independent set of size $\alpha+1$ in $Sp$, which is absurd. ∎
Two vertices $u,v$ in a graph $G$ are said to be twins in $G$ if
$N_{G}(u)\setminus v=N_{G}(v)\setminus u.$
It is easy to check that “to be twins” is an equivalence relation over $V(G)$, and the equivalence classes are cliques or independent sets of $G$.
Theorem 4.4.
If $Sp$ is an unbalanced split graph, then
$A_{Sp}=W(Sp),\quad B_{Sp}=K^{*}(Sp),\quad C_{Sp}=S^{*}(Sp).$
Moreover, $W(Sp)$ is always a clique or an independent set, of mutually twin vertices.
Proof.
Since $A\dot{\cup}B\dot{\cup}C=V(Sp)=W\dot{\cup}K^{*}\dot{\cup}S^{*}$, the required identities follows from Sections 4, 4 and 4. The statement about $W$ can be subsequently derived from Theorem 2 in [3], using that $A=W$. ∎
Corollary 4.5.
Let $Sp(K,S)$ be a connected split graph such that $K=\bigcup_{v\in S}N_{Sp}(v)$ and $\operatorname{rk}(R)=1$. Then:
@@
    $R=J$;
@@
    all the independent (clique) vertices of $Sp$ are pairwise twins;
@@
    $(K,S)=(K^{*}(Sp),W(Sp))$ and $(|K|,|S|)=(\omega(Sp)-1,\alpha(Sp))$.
Proof.
  1. (1).
    Since $Sp$ is connected and has no isolated vertices in $S$, all the columns of $R$ are nonzero. Since $K=\bigcup_{v\in S}N(v)$, all the rows of $R$ are nonzero. Moreover, since $\operatorname{rk}(R)=1$, all the columns of $R$ must be scalar multiples of each other. Since all columns are nonzero $\{0,1\}$-vectors, they must all be equal to the same vector. Since $R_{i*}\neq 0$ for all $i$ and $R$ is a $\{0,1\}$-matrix, we get that $R_{*j}=\operatorname{1\!\!1}$ for all $j$.
  2. (2).
    Since $R=J$, we have $N(v)=K$ for all $v\in S$ and $N(u)=(K\setminus u)\cup S$ for all $u\in K$.
  3. (3).
    If $|K|=\omega$ and $v\in S$, then $K\cup v$ would be a clique of size $\omega+1$ in $Sp$, which is absurd. Then, $\deg(v)=\omega-1$, and hence $S\subset A=W$ by Theorem 4.4. On the other hand, since every clique vertex of $Sp$ is universal, we have $K\subset B=K^{*}$ by Theorem 4.4. ∎
Proposition 4.6.
@@
    Let $K$ be a maximum clique of $Sp$. Since $Sp$ is unbalanced, we know that $\alpha+\omega=|Sp|+1$, and so $\alpha-1=|Sp|-\omega=|V(Sp)\setminus K|$. Let $S$ be a maximum independent set of $Sp$. Since $|K\cap S|\leq 1,|S|=\alpha$ and $|V(Sp)\setminus K|=\alpha-1$, it must be $S=(V(Sp)\setminus K)\dot{\cup}w$, for some $w\in K$. Consequently, $V(Sp)\setminus K$ is an independent set of $Sp$. Thus, $(K,V(Sp)\setminus K)$ is an s-partition for $Sp$, which means that $K\in\mathcal{K}(Sp)$.
@@
    Let $S$ be a maximum independent set of $Sp$. Then, $S$ is a maximum clique of the graph $\overline{Sp}$ (the complement of $Sp$), which is clearly split and unbalanced. From (1), we know that $S\in\mathcal{K}(\overline{Sp})$, which means that $V(\overline{Sp})\setminus S$ is an independent set of $\overline{Sp}$. Hence, $V(Sp)\setminus S$ is a clique of $Sp$, which shows that $S\in\mathcal{S}(Sp)$. ∎
Section 4 does not hold in general for balanced split graph. In fact, every balanced $Sp(K,S)$ with an independent vertex of degree |K|1|K|-1 is a counterexample ($P_{4}$ is the “simplest” one). In other words, a balanced split graph $Sp(K,S)$ has always a unique s-partition, but $K$ might not be the only maximum clique of $Sp$. An analogous phenomenon can be observed for maximum independent sets. The following result clarifies this behavior of the balanced split graphs.
@@
    If $x$ is a vertex of $Sp$ of degree |K|1|K|-1, then $x$ is an independent vertex of $G$ (otherwise, $x$ would be swing). Now, observe that $N(x)\cup x$ is a maximum clique of $Sp$, different from $K$.
    For the converse, suppose that $K^{\prime}$ is a maximum clique of $Sp$, different from $K$. Since $|K^{\prime}\cap S|\leq 1$, it must be $K^{\prime}=(K\setminus w)\dot{\cup}x$, for some $w\in K$ and $x\in S$. Since $x\in K^{\prime}$, we have $K\setminus w\subset N(x)$. Since $\in S, xx$ is not adjacent to any vertex of $S$, which means that $N(x)\subset K$. If $w\in N(x)$, then $N(x)=K$ and hence $x\in W(Sp)=\varnothing$. This absurdity implies $N(x)=K\setminus w$, i.e., $\deg(x)=|K|-1$.
@@
    If $x$ is a vertex of $Sp$ such that $N(x)\cap S=\{w\}$ for some $w$, it is clear that $x\in K$ and $(I\setminus w)\dot{\cup}x$ is a maximum independent set different from $S$.
    For the converse, suppose that $S^{\prime}$ is a maximum independent set of $Sp$, different from $S$. Since $|K\cap S^{\prime}|\leq 1$, it must be $S^{\prime}=(S\setminus w)\dot{\cup}x$, for some $w\in S$ and $x\in K$. Since $x\in K$, $K\setminus x\subset N(x)$. Since $\in S^{\prime}, xx$ is not adjacent to any vertex of $S\setminus w$. Notice that $w\in N(x)$ (otherwise, $x$ would be swing). Therefore, $N(x)=(K\setminus x)\dot{\cup}w$, which means that $|N(x)\cap S|=1$. ∎
Corollary 4.8.
  1. 1.
    If $W=\{w\}$, then
    $\mathcal{KS}=\{(K^{*}\cup w,S^{*}),(K^{*},S^{*}\cup w)\};$
    moreover, $K^{*}\cup w$ is the unique maximum clique of $Sp$ and $S^{*}\cup w$ is the unique maximum independent set of $Sp$.
  2. 2.
    If $W$ is a clique of size $\geq 2$, then
    $\mathcal{KS}=\{(K^{*}\cup W,S^{*})\}\cup\{(K^{*}\cup(W\setminus w),S^{*}\cup w):w\in W\};$
    moreover, $K^{*}\cup W$ is the unique maximum clique of $Sp$, and all the maximum independent sets of $Sp$ are of the form $S^{*}\cup w$.
  3. 3.
    If $W$ is an independent set of size $\geq 2$, then
    $\mathcal{KS}=\{(K^{*},S^{*}\cup W)\}\cup\{(K^{*}\cup w,S^{*}\cup(W\setminus w)):w\in W\};$
    moreover, $S^{*}\cup W$ is the unique maximum independent set of $Sp$, and all the maximum cliques of $Sp$ are of the form $K^{*}\cup w$.
Proof.
It follows from Theorem 4.4, Theorem 10 in [3] and Section 4. ∎
Notice that Section 4 implicitly characterizes a swing vertex in terms of its neighborhood (compare with (2)):
  1. (1).
    If $W$ is a clique, then $w$ is swing if and only if $N(w)=K^{*}\cup(W\setminus w)=\{v\neq w:\deg(v)\geq\omega-1\}$;
  2. (2).
    If $W$ is an independent set, then $w$ is swing if and only if $N(w)=K^{*}=\{v:\deg(v)\geq\omega\}$.
The following theorem states that in unbalanced split graphs, supported vertices must lie in the intersection of all the maximum independent sets.
@@
Consider an arbitrary s-partition $(K,S)$ for $Sp$ such that $|S|=\alpha(Sp)$. The existence of such is guaranteed by Section 4. Then, by Section 4, there exists a vertex $w\in S$ such that $N(w)=K$, which means that $R_{*w}=\operatorname{1\!\!1}$. Since $\operatorname{1\!\!1}\in\operatorname{\mathcal{R}}(R)$, we conclude by Theorem 3.2(5) that $\operatorname{Supp}(Sp)\subset S$. Therefore, $\operatorname{Supp}(Sp)$ is contained in the intersection of all the maximum independent sets in $\mathcal{S}(Sp)$. Now, use Section 4(2). ∎
@@
If $G$ is a threshold graph, then $\operatorname{Supp}(G)$ is contained in the intersection of all the maximum independent sets of $G$.
@@
It follows from the previous discussion and from Theorem 4.9. ∎